OpenStudy (anonymous):

Read the proof. Fill in the empty justifications. Conjecture: If 3(x - 1) = -27, then x = -8. 3(x - 1) = -27 A. Given 3x - 3 = -27 B. _________________ 3x = -24 C. _________________ x = -8 D. _________________

OpenStudy (jiteshmeghwal9):

Take x=-8 in each equation. for example-> 3(x-1)=-27=>3(-8-1)=27H.P.

OpenStudy (anonymous):

im still confused.

OpenStudy (jiteshmeghwal9):

3x-3=-27 put -8 instead of x 3(-8)-3=-27 now solve it if u gt both sides equal then it's right otherwise it's wrong.

OpenStudy (anonymous):

oh okay , thank you bunches!

OpenStudy (jiteshmeghwal9):

yw :)